OpenStudy (anonymous):

HELP! 2t+8 is greater than or equal to -4(t+1) t is greater than or equal to?

OpenStudy (anonymous):

this might make more sense \[2t+8\ge-4(t-1)\]

OpenStudy (anonymous):

so \[t \ge?\]

OpenStudy (antonio_xx2):

i think greater than

TheSmartOne (thesmartone):

Distribute the -4 \[2t+8\ge-4t+4\] Move like terms to one side \[6t \ge -4\] Divide by 6 to get t alone \[t \ge \frac{ -2 }{ 3 }\]